# From Competition to Coordination: Market Making as a Scalable Framework for Safe and Aligned Multi-Agent LLM Systems

### v1

@@ 1 Introduction
The rapid deployment of artificial intelligence systems across safety-critical domains has intensified concerns regarding existential risks, particularly the emergence of deceptively aligned models (hendrycks2023overviewcatastrophicairisks). Recent evidence demonstrates that advanced language models exhibit strategic deception, including attempts to game evaluation protocols and misrepresent their internal states during training (carlsmith2023schemingaisaisfake). These alignment failures manifest as sycophancy, systematic untruthfulness, and adversarial behaviour, actions that empirically worsen with increased model scale (ji2025aialignmentcomprehensivesurvey).
Existing alignment methodologies face fundamental limitations. Reinforcement Learning from Human Feedback (RLHF), while effective for surface-level behavioural modification, remains vulnerable to reward hacking and evaluator deception. Debate-based approaches require human adjudication that cannot scale to superhuman reasoning capabilities. The Alignment Research Center’s Eliciting Latent Knowledge (ELK) framework defines this challenge: extracting model’s true internal representations rather than their strategically chosen outputs (christiano2022elksummary).
@@ 2.1 The Challenges of Control
amodei2016concreteproblemsaisafety established a taxonomy of AI safety failure modes comprising five critical categories: negative side effects, reward hacking, scalable oversight limitations, unsafe exploration, and distributional shift vulnerabilities. This foundational framework reveals an inherent tension in alignment objectives: excessive optimization for harmlessness produces ineffectual systems, while prioritizing capability enables potential misuse (bai2022traininghelpfulharmlessassistant). The multidimensional nature of these constraints implies that no single methodology can simultaneously address all failure modes, necessitating approaches that optimize across multiple safety dimensions.
@@ 2.1.1 Human-Centric Alignment
Early alignment techniques relied on direct human supervision through iterative feedback mechanisms. Reinforcement Learning from Human Feedback (RLHF) exemplifies this paradigm, wherein human evaluators shape model behaviour through preference rankings (bai2022traininghelpfulharmlessassistant). AI Safety via Debate was proposed by irving2018aisafetydebate, structuring oversight as adversarial argumentation adjudicated by human judges. These approaches face three fundamental limitations. First, the bandwidth constraint: human evaluation cannot scale to the volume and velocity of decisions required in deployed systems (amodei2016concreteproblemsaisafety). Second, the competence boundary: superhuman AI capabilities exceed human evaluators’ ability to assess correctness (ji2025aialignmentcomprehensivesurvey). Third, the alignment targeting problem, where models optimize for evaluator approval rather than ground truth, leading to sycophantic behaviour and strategic deception (carlsmith2023schemingaisaisfake; park2023aideceptionsurveyexamples).
@@ 2.1.2 AI-Mediated Oversight
These challenges of human-centric alignment motivated the use of AI-mediated oversight, in which secondary AI systems help supervise the model being tested, aiming to augment or replace human adjudicators. JudgeLM, replacing the human judge in AI debate with an AI system, demonstrates extended capabilities in a variety of situations (zhu2025judgelmfinetunedlargelanguage). bowman2022measuringprogressscalableoversight also explore how less-capable AIs can reliably evaluate stronger ones without expert human intervention (bowman2022measuringprogressscalableoversight). Together, these efforts suggest that scalable, AI-mediated oversight may be a necessary step toward maintaining safe and reliable control as AI capabilities continue to grow.
@@ 2.2 Market Making as a Control Mechanism
Market making offers an incentive-based alternative to adjudication. In this method, an automated market maker posts prices for propositions and traders (models or submodels) buy or sell claims based on their beliefs; iterative trading drives prices toward an equilibrium that reflects collective credence (holmes2020aisafetymarketmaking). A key intended advantage is enforcing myopic behavior where trader agents optimize per-step trades, reducing incentives for long-term scheming that can undermine debate or RLHF. Market mechanisms also facilitate per-step inspection, probabilistic scoring, and potential scalability without continuous human adjudication. Thus, market making attempts to improve upon existing methods of AI control.
Practical challenges to market making do remain, including defending the market maker against false claims, designing proper rewards that prevent gaming, and ensuring robust performance when truth is hazy. Existing work is primarily a toy implementation of market making from Cameron Holmes holmes2020aisafetymarketmaking.
@@ 4.1 Model Selection
GPT Model Family. We tested five variants from the GPT family: GPT-4.1 nano, GPT-4.1 mini, GPT-4.1, gpt-oss-20b, and gpt-oss-120b (openai2024gpt4ocard; openai2025gptoss120bgptoss20bmodel). This selection spans from lightweight models (nano, mini) to large-scale models with enhanced reasoning capabilities. The inclusion of both proprietary (GPT-4.1 series) and open-source variants allows us to better assess whether market making effectiveness depends on specific training methodologies or remains consistent across development paradigms.
Qwen3 Model Family. The Qwen3 (yang2025qwen3technicalreport) family provided our most comprehensive scale analysis, with seven model sizes: 0.6B, 1.7B, 4B, 8B, 14B, 32B, and 235B parameters. This extensive range enables us to examine how market making behaviour scales with model capacity, particularly whether larger models exhibit more sophisticated trading strategies or demonstrate increased susceptibility to adversarial arguments.
Llama 3 Model Family. We evaluated three Llama 3 (grattafiori2024llama3herdmodels) variants (3B, 8B, and 70B parameters) to assess market making performance scaling across model capacity and training objectives.
@@ 4.2 Dataset Selection
TruthfulQA. TruthfulQA (lin2022truthfulqameasuringmodelsmimic) serves as our primary benchmark for factual accuracy, containing questions specifically designed to elicit false beliefs or misconceptions that models may have learned from training data. In the market making context, this dataset tests whether iterative trading can correct initial false predictions, with the trader model potentially identifying and challenging spurious correlations or misconceptions held by the market maker.
Scruples (Dilemmas). The Scruples (lourie2021scruplescorpuscommunityethical) dataset, specifically the Dilemmas subset presents real-world ethical dilemmas sourced from online advice forums, requiring models to reason about complex moral scenarios without clear-cut answers. With regards to market making, Scruples tests whether the trading mechanism can navigate moral ambiguity and converge on socially acceptable judgments in everyday ethical situations.
ETHICS (Justice, Commonsense). We select two subsets from the ETHICS (hendrycks2023aligningaisharedhuman) dataset to evaluate different aspects of moral reasoning. The Justice subset tests understanding of fairness and impartiality, central to many alignment objectives. The Commonsense subset evaluates basic moral intuitions that should be robust across cultural contexts.
CommonsenseQA 2.0. CommonsenseQA 2.0 (talmor2022commonsenseqa20exposinglimits) provides a test of general reasoning and world knowledge, requiring models to make inferences based on everyday situations. As opposed to the original CommonsenseQA, the 2.0 version includes adversarially-filtered questions that challenge models’ reasoning capabilities.
@@ 5 Results
The Qwen family of models had the highest overall increase in percentage accuracy for each dataset, reaching an increase of almost $10\%$ in ETHICS-J and over $5\%$ across the board. These gains are significant improvements over the models’ individual baselines. One explanation for this is Qwen3’s extensive post-training pipeline emphasizing chain-of-thought reasoning as a core architectural feature alongside its unified thinking and non-thinking architecture (yang2025qwen3technicalreport).

### v2

@@ 1 Introduction
The rapid deployment of artificial intelligence systems across safety-critical domains has intensified concerns regarding existential risks, particularly the emergence of deceptively aligned models (Hendrycks et al., 2023b). Recent evidence demonstrates that advanced language models exhibit strategic deception, including attempts to game evaluation protocols and misrepresent their internal states during training (Carlsmith, 2023). These alignment failures manifest as sycophancy, systematic untruthfulness, and adversarial behaviour, actions that empirically worsen with increased model scale (Ji et al., 2025).
Existing alignment methodologies face fundamental limitations. Reinforcement Learning from Human Feedback (RLHF), while effective for surface-level behavioural modification, remains vulnerable to reward hacking and evaluator deception. Debate-based approaches require human adjudication that cannot scale to superhuman reasoning capabilities. The Alignment Research Center’s Eliciting Latent Knowledge (ELK) framework defines this challenge: extracting model’s true internal representations rather than their strategically chosen outputs (Christiano et al., 2022).
@@ 2.1 The Challenges of Control
Amodei et al. (2016) established a taxonomy of AI safety failure modes comprising five critical categories: negative side effects, reward hacking, scalable oversight limitations, unsafe exploration, and distributional shift vulnerabilities. This foundational framework reveals an inherent tension in alignment objectives: excessive optimization for harmlessness produces ineffectual systems, while prioritizing capability enables potential misuse (Bai et al., 2022). The multidimensional nature of these constraints implies that no single methodology can simultaneously address all failure modes, necessitating approaches that optimize across multiple safety dimensions.
@@ 2.1.1 Human-Centric Alignment
Early alignment techniques relied on direct human supervision through iterative feedback mechanisms. Reinforcement Learning from Human Feedback (RLHF) exemplifies this paradigm, wherein human evaluators shape model behaviour through preference rankings (Bai et al., 2022). AI Safety via Debate was proposed by Irving et al. (2018), structuring oversight as adversarial argumentation adjudicated by human judges. These approaches face three fundamental limitations. First, the bandwidth constraint: human evaluation cannot scale to the volume and velocity of decisions required in deployed systems (Amodei et al., 2016). Second, the competence boundary: superhuman AI capabilities exceed human evaluators’ ability to assess correctness (Ji et al., 2025). Third, the alignment targeting problem, where models optimize for evaluator approval rather than ground truth, leading to sycophantic behaviour and strategic deception (Carlsmith, 2023; Park et al., 2023).
@@ 2.1.2 AI-Mediated Oversight
These challenges of human-centric alignment motivated the use of AI-mediated oversight, in which secondary AI systems help supervise the model being tested, aiming to augment or replace human adjudicators. JudgeLM, replacing the human judge in AI debate with an AI system, demonstrates extended capabilities in a variety of situations (Zhu et al., 2025). Bowman et al. (2022) also explore how less-capable AIs can reliably evaluate stronger ones without expert human intervention (Bowman et al., 2022). Together, these efforts suggest that scalable, AI-mediated oversight may be a necessary step toward maintaining safe and reliable control as AI capabilities continue to grow.
@@ 2.2 Market Making as a Control Mechanism
Market making offers an incentive-based alternative to adjudication. In this method, an automated market maker posts prices for propositions and traders (models or submodels) buy or sell claims based on their beliefs; iterative trading drives prices toward an equilibrium that reflects collective credence (Holmes, 2020). A key intended advantage is enforcing myopic behavior where trader agents optimize per-step trades, reducing incentives for long-term scheming that can undermine debate or RLHF. Market mechanisms also facilitate per-step inspection, probabilistic scoring, and potential scalability without continuous human adjudication. Thus, market making attempts to improve upon existing methods of AI control.
Practical challenges to market making do remain, including defending the market maker against false claims, designing proper rewards that prevent gaming, and ensuring robust performance when truth is hazy. Existing work is primarily a toy implementation of market making from Cameron Holmes Holmes (2020).
@@ 4.1 Model Selection
GPT Model Family. We tested five variants from the GPT family: GPT-4.1 nano, GPT-4.1 mini, GPT-4.1, gpt-oss-20b, and gpt-oss-120b (OpenAI et al., 2024, 2025). This selection spans from lightweight models (nano, mini) to large-scale models with enhanced reasoning capabilities. The inclusion of both proprietary (GPT-4.1 series) and open-source variants allows us to better assess whether market making effectiveness depends on specific training methodologies or remains consistent across development paradigms.
Qwen3 Model Family. The Qwen3 (Yang et al., 2025) family provided our most comprehensive scale analysis, with seven model sizes: 0.6B, 1.7B, 4B, 8B, 14B, 32B, and 235B parameters. This extensive range enables us to examine how market making behaviour scales with model capacity, particularly whether larger models exhibit more sophisticated trading strategies or demonstrate increased susceptibility to adversarial arguments.
Llama 3 Model Family. We evaluated three Llama 3 (Grattafiori et al., 2024) variants (3B, 8B, and 70B parameters) to assess market making performance scaling across model capacity and training objectives.
@@ 4.2 Dataset Selection
TruthfulQA. TruthfulQA (Lin et al., 2022) serves as our primary benchmark for factual accuracy, containing questions specifically designed to elicit false beliefs or misconceptions that models may have learned from training data. In the market making context, this dataset tests whether iterative trading can correct initial false predictions, with the trader model potentially identifying and challenging spurious correlations or misconceptions held by the market maker.
Scruples (Dilemmas). The Scruples (Lourie et al., 2021) dataset, specifically the Dilemmas subset presents real-world ethical dilemmas sourced from online advice forums, requiring models to reason about complex moral scenarios without clear-cut answers. With regards to market making, Scruples tests whether the trading mechanism can navigate moral ambiguity and converge on socially acceptable judgments in everyday ethical situations.
ETHICS (Justice, Commonsense). We select two subsets from the ETHICS (Hendrycks et al., 2023a) dataset to evaluate different aspects of moral reasoning. The Justice subset tests understanding of fairness and impartiality, central to many alignment objectives. The Commonsense subset evaluates basic moral intuitions that should be robust across cultural contexts.
CommonsenseQA 2.0. CommonsenseQA 2.0 (Talmor et al., 2022) provides a test of general reasoning and world knowledge, requiring models to make inferences based on everyday situations. As opposed to the original CommonsenseQA, the 2.0 version includes adversarially-filtered questions that challenge models’ reasoning capabilities.
@@ 5 Results
The Qwen family of models had the highest overall increase in percentage accuracy for each dataset, reaching an increase of almost $10\%$ in ETHICS-J and over $5\%$ across the board. These gains are significant improvements over the models’ individual baselines. One explanation for this is Qwen3’s extensive post-training pipeline emphasizing chain-of-thought reasoning as a core architectural feature alongside its unified thinking and non-thinking architecture (Yang et al., 2025).